OpenStudy (anonymous):

what is -1/2x+2=-x+7

OpenStudy (anonymous):

equation

OpenStudy (anonymous):

you could add -1/2x to both side. which would be what? you tell me.

OpenStudy (anonymous):

1 and 1/2

OpenStudy (anonymous):

right

OpenStudy (anonymous):

yes so then the equation would be then be 1/2x+2=7

OpenStudy (anonymous):

can do the rest or do still need more help?

OpenStudy (anonymous):

1 and 1/2x=5

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so you good??

OpenStudy (anonymous):

no im stuck there

OpenStudy (anonymous):

so you got 1/2x = 5 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is it 10

OpenStudy (anonymous):

yes!!!

OpenStudy (anonymous):

you got it ;)

OpenStudy (anonymous):

can you help me with another

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

5(x + 8) = 7(x - 2)

OpenStudy (anonymous):

i got to 5x+40=7x-14

OpenStudy (anonymous):

you could distribute the equation first

OpenStudy (anonymous):

then subtract 5x for 7x

OpenStudy (anonymous):

-2x

OpenStudy (anonymous):

right, then the equation would be -2x + 40 = -14

OpenStudy (anonymous):

would i minus the 14 from 40 or add 40 to 14

OpenStudy (anonymous):

right, now go on

OpenStudy (anonymous):

which one do i do add or subtract

OpenStudy (anonymous):

anyone you like, just pick

OpenStudy (anonymous):

im stuck at -2x+40=-14

OpenStudy (anonymous):

-2x=-54

OpenStudy (anonymous):

27

OpenStudy (anonymous):

just to make it easier minus 40 from -14 to get -54, then the equation would be; 2x=-54

OpenStudy (anonymous):

-(1/2x)+2=-x+7 so add 1/2 to both sides 2 = -x + (1/2x) +7 combine the -x and the 1/2 -5 = (-1/2x) devide both sides by (-1/2) and finally x = 10

OpenStudy (anonymous):

i meant -2x=-54

OpenStudy (anonymous):

4 - (x + 2) = 6 - 5(x + 4)

OpenStudy (anonymous):

which would be 27 like you got :)

OpenStudy (anonymous):

so like with the one before distribute first

OpenStudy (anonymous):

idk what you guy's are doing -1 + .5 = -.5

OpenStudy (anonymous):

which be 4 - x-2=6-5x-20

OpenStudy (anonymous):

in the first one

OpenStudy (anonymous):

Huh??

OpenStudy (anonymous):

-(1/2x)+2=-x+7 so add 1/2 to both sides 2 = -x + (1/2x) +7 combine the -x and the 1/2 -5 = (-1/2x) devide both sides by (-1/2) and finally x = 10

OpenStudy (anonymous):

yeay what about it??

OpenStudy (anonymous):

thats what we got

OpenStudy (anonymous):

fuk sry I thought you got -54

OpenStudy (anonymous):

NVM

OpenStudy (anonymous):

@cierajeanhodges can you do the rest

OpenStudy (anonymous):

ANYWAY IM STILL STUCK WITH THE PROBLEM WE ARE WORKING

OpenStudy (anonymous):

ok so what do not get?

OpenStudy (anonymous):

all of it

OpenStudy (anonymous):

so can you distribute the equation first

OpenStudy (anonymous):

i did what you did

OpenStudy (anonymous):

so what did you get?

OpenStudy (anonymous):

4-x-2=6-5x-20 and im stuck

OpenStudy (anonymous):

so now try to get all the numbers on one side and x on the other side

OpenStudy (anonymous):

so the equation like this: 4x=-16

OpenStudy (anonymous):

i'm sure you can do the rest ;)

OpenStudy (anonymous):

so need anymore help???

OpenStudy (anonymous):

@cierajeanhodges you there?